Not sure how to solve this equation.
5x + 8 + 3x - x + 5 = 6x -3
Thanks
Kathy

Answer by checkley75(3666)

5x+8+3x-x+5=6x-3
5x+3x-x-6x=-3-8-5
x=-16
proof
5*-16+8+3*-16+16+5=6*-16-3
-80+8-48-16+5=-96-3
-99=-99